# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 1 Introduction and Main Results
For clarity, we present the proof in three steps.
Step 1. Fix $\varepsilon>0$ and define the functions:
$\theta_{\varepsilon}=e^{\lambda\alpha_{\varepsilon}},\qquad\alpha_{\varepsilon}\equiv\alpha_{\varepsilon}(t,x)=\frac{e^{\mu\psi(x)}-e^{2\mu|\psi|_{\infty}}}{(t+\varepsilon)(T-t+\varepsilon)}.$ (2.4)
$\inf\big\{J_{\varepsilon}(u,v_{1},v_{2})\,|\,(u,v_{1},v_{2})\in\mathcal{U}\big\},$ (2.5)
where
$J_{\varepsilon}(u,v_{1},v_{2})=$ $\,\frac{1}{2}\mathbb{E}\iint_{Q_{0}}\lambda^{-3}\mu^{-4}\theta^{-2}\varphi^{-3}u^{2}dxdt+\frac{1}{2}\mathbb{E}\iint_{Q}\lambda^{-2}\mu^{-2}\theta^{-2}\varphi^{-2}v_{1}^{2}dxdt$
$+\frac{1}{2}\mathbb{E}\iint_{\Sigma}\lambda^{-2}\mu^{-2}\theta^{-2}\varphi^{-2}v_{2}^{2}d\sigma dt+\frac{1}{2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y^{2}dxdt+\frac{1}{2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\Gamma}^{2}d\sigma dt$
and
$\mathcal{U}=\Bigg\{($ $u,v_{1},v_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)):$
$\mathbb{E}\iint_{Q_{0}}\theta^{-2}\varphi^{-3}u^{2}dxdt<\infty\,,\quad\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{1}^{2}dxdt<\infty\,,\quad\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{2}^{2}d\sigma dt<\infty\Bigg\}.$
$u_{\varepsilon}=-\mathbbm{1}_{G_{0}}\lambda^{3}\mu^{4}\theta^{2}\varphi^{3}r_{\varepsilon}\,\,,\quad\quad v_{\varepsilon,1}=-\lambda^{2}\mu^{2}\theta^{2}\varphi^{2}R_{\varepsilon,1}\,\,,\quad\quad v_{\varepsilon,2}=-\lambda^{2}\mu^{2}\theta^{2}\varphi^{2}R_{\varepsilon,2},$ (2.6)
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &=\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}zr_{\varepsilon}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}r_{\varepsilon,\Gamma}d\sigma dt.\end{aligned}$ (2.8)
By Young’s inequality, we get from (2.8) that for all $\rho>0$,
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &\leq\rho\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}r_{\varepsilon,\Gamma}^{2}d\sigma\Bigg]\\ &\;\;\;+\frac{1}{4\rho}\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$ (2.9)
Using the Carleman estimate (2.1) for solutions of (2.7) and noting that $\theta\theta_{\varepsilon}^{-1}\leq 1$, then there exists a constant $C=C(G,G_{0},\mathcal{A},\mathcal{A}_{\Gamma})>0$ such that for all $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, the inequality (2.9) implies that
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\rho\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &\qquad\;\quad+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\Bigg]\\ &\;\;\;+\frac{1}{4\rho}\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$ (2.10)
Choosing a small enough $\rho>0$ in (2.10), we obtain that
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$
Recalling (2.6), it follows that
 $\begin{aligned} &\,2\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}\mathcal{A}\nabla y_{\varepsilon}\cdot\nabla y_{\varepsilon}dxdt\\ &=-2\mathbb{E}\iint_{Q}y_{\varepsilon}\mathcal{A}\nabla y_{\varepsilon}\cdot\nabla(\theta_{\varepsilon}^{-2}\varphi^{-2})dxdt+2\mathbb{E}\int_{0}^{T}\langle\partial_{\nu}^{\mathcal{A}}y_{\varepsilon},\theta_{\varepsilon}^{-2}\varphi^{-2}y_{\varepsilon,\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt\\ &\quad\,+\mathbb{E}\iint_{Q}(\theta^{-2}_{\varepsilon}\varphi^{-2})_{t}y_{\varepsilon}^{2}dxdt+2\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\theta^{2}\varphi y_{\varepsilon}zdxdt\\ &\quad\,+2\mathbb{E}\iint_{Q_{0}}\theta^{-2}_{\varepsilon}\varphi^{-2}y_{\varepsilon}u_{\varepsilon}dxdt+\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt.\end{aligned}$ (2.12)
$|(\theta^{-2}_{\varepsilon}\varphi^{-2})_{t}|\leq CT\lambda e^{2\mu|\psi|_{\infty}}\theta_{\varepsilon}^{-2},\qquad|\nabla(\theta^{-2}_{\varepsilon}\varphi^{-2})|\leq C\lambda\mu\theta_{\varepsilon}^{-2}\varphi^{-1}.$ (2.14)
 $\begin{aligned} &\,2\beta_{0}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda\mu\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}\varphi^{-1}|y_{\varepsilon}||\nabla y_{\varepsilon}|dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+2\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon}||z|dxdt\\ &\quad\,+2\mathbb{E}\iint_{Q_{0}}\theta^{-1}_{\varepsilon}\theta^{-1}\varphi^{-2}|y_{\varepsilon}||u_{\varepsilon}|dxdt+\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &\quad\,+2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon,\Gamma}||z_{\Gamma}|d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.15)
 $\begin{aligned} &\,2\beta_{0}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq\rho\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\frac{C}{\rho}\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt\\ &\quad\,+2\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon}||z|dxdt+2\mathbb{E}\iint_{Q_{0}}\theta^{-1}_{\varepsilon}\theta^{-1}\varphi^{-2}|y_{\varepsilon}||u_{\varepsilon}|dxdt+\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt\\ &\quad\,+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon,\Gamma}||z_{\Gamma}|d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.16)
$\begin{aligned} &\,\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon}||z|dxdt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q_{0}}\theta^{-1}_{\varepsilon}\theta^{-1}\varphi^{-2}|y_{\varepsilon}||u_{\varepsilon}|dxdt+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt\\ &\quad\,+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon,\Gamma}||z_{\Gamma}|d\sigma dt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.17)
Using Young’s inequality for (2.17), we have that
$\begin{aligned} &\,\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+C\lambda^{2}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}dxdt\\ &\quad\,+C\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+C\lambda^{-4}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-4}u_{\varepsilon}^{2}dxdt+C\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$
Taking a large enough $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$ and $\mu\geq C$, it follows that
Combining (2.18) and (2.11), we end up with
which implies that
Then, we deduce that there exists
$(\widehat{u},\widehat{v}_{1},\widehat{v}_{2},\widehat{y},\widehat{y}_{\Gamma})\in$ $\,\,L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$
such that as $\varepsilon\rightarrow 0$,
$u_{\varepsilon}\longrightarrow\widehat{u},\,\,\,\textnormal{weakly in}\,\,\,L^{2}((0,T)\times\Omega;L^{2}(G_{0}));$
@@
$\begin{cases}\begin{array}[]{ll}d\phi+\nabla\cdot(\mathcal{A}\nabla\phi)\,dt=f\,dt+\Phi\,dW(t)&\textnormal{in}\,\,Q,\\ d\phi_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}\phi_{\Gamma})dt-\partial_{\nu}^{\mathcal{A}}\phi\,dt=g\,dt+\widehat{\Phi}\,dW(t)&\textnormal{on}\,\,\Sigma,\\ \phi_{\Gamma}(t,x)=\phi|_{\Gamma}(t,x)&\textnormal{on}\,\,\Sigma,\\ (\phi,\phi_{\Gamma})|_{t=T}=(0,0)&\textnormal{in}\,\,G\times\Gamma.\end{array}\end{cases}$
Letting $\varepsilon\rightarrow 0$ in (2.21), we get
Now, we are in a position to prove Theorem 1.1.
Proof of Theorem 1.1.

### v2

@@ 2 Preliminaries and Well-Posedness Results
For the uniqueness, since the system is linear, we assume that
$(z,z_{\Gamma};Z,\widehat{Z})\in\Big(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\Big)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})$
is a weak solution to (1.3) with vanishing data, that is, with terminal condition $(z_{T},z_{\Gamma,T})=(0,0)$ and source terms $F_{1}=F_{2}=F=F_{\Gamma}=0$. Fix $t\in[0,T]$, applying Itô’s formula as stated in Lemma 2.2 to $\|(z,z_{\Gamma})\|_{\mathbb{L}^{2}}^{2}$, computing $d_{s}\|(z,z_{\Gamma})\|_{\mathbb{L}^{2}}^{2}$, and integrating over the interval $s\in(t,T]$, we obtain
$\mathbb{E}\int_{G}|z(t)|^{2}\,dx+\mathbb{E}\int_{\Gamma}|z_{\Gamma}(t)|^{2}\,d\sigma$
$=-2\mathbb{E}\int_{t}^{T}\int_{G}\mathcal{A}\nabla z\cdot\nabla z\,dx\,ds-2\mathbb{E}\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma\,ds$
$\quad-\mathbb{E}\int_{t}^{T}\int_{G}|Z|^{2}\,dx\,ds-\mathbb{E}\int_{t}^{T}\int_{\Gamma}|\widehat{Z}|^{2}\,d\sigma\,ds.$
Recalling assumption (1.1), it follows that
$\mathbb{E}\int_{G}|z(t)|^{2}\,dx+\mathbb{E}\int_{\Gamma}|z_{\Gamma}(t)|^{2}\,d\sigma+\mathbb{E}\int_{t}^{T}\int_{G}|Z|^{2}\,dx\,ds+\mathbb{E}\int_{t}^{T}\int_{\Gamma}|\widehat{Z}|^{2}\,d\sigma\,ds$
$\leq-2\beta_{0}\mathbb{E}\int_{t}^{T}\int_{G}|\nabla z|^{2}\,dx\,ds-2\beta_{0}\mathbb{E}\int_{t}^{T}\int_{\Gamma}|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma\,ds$
$\leq 0.$
Consequently, we conclude that
$(z,z_{\Gamma};Z,\widehat{Z})\equiv(0,0;0,0),\quad\textnormal{a.s.}$
This completes the proof of the uniqueness of solutions to system (1.3).
For the existence, the approach consists of decomposing the solution into two parts: the solution of an elliptic problem and the solution of a backward stochastic parabolic equation without the divergence terms, for which well-posedness is already known from Proposition 2.2. Subsequently, by employing a sequential argument, we establish the existence of a solution to (1.2). To this end, let $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, $F_{1}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, $F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, and assume that $(F,F_{\Gamma})\in\mathscr{F}$, where
$\mathscr{F}=\Big\{$ $(F,F_{\Gamma})\in L^{2}_{\mathcal{F}}(\Omega;C^{1}([0,T];L^{2}(G;\mathbb{R}^{N})))\times L^{2}_{\mathcal{F}}(\Omega;C^{1}([0,T];L^{2}(\Gamma;\mathbb{R}^{N}))):$
$\quad(F(T,\cdot),F_{\Gamma}(T,\cdot))=(0,0)\;\;\textnormal{in}\;G\times\Gamma,\;\textnormal{a.s.}\Big\},$
which is dense in $L^{2}_{\mathcal{F}}(0,T;L^{2}(G;\mathbb{R}^{N}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma;\mathbb{R}^{N}))$. Hence, the elliptic system
@@
$\begin{cases}\nabla\cdot(\mathcal{A}\nabla h)=\nabla\cdot F&\text{in }Q,\\[2.84526pt] \nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})-\partial_{\nu}^{\mathcal{A}}h=\nabla_{\Gamma}\cdot F_{\Gamma}-F\cdot\nu&\text{on }\Sigma,\\[2.84526pt] h_{\Gamma}=h|_{\Gamma}&\text{on }\Sigma,\end{cases}$ (2.4)
admits a unique solution $(h,h_{\Gamma})\in L^{2}_{\mathcal{F}}(\Omega;C^{1}([0,T];\mathbb{H}^{1}))$ such that the terminal condition
$(h(T,\cdot),h_{\Gamma}(T,\cdot))=(0,0)\quad\text{in }G\times\Gamma,\;\text{a.s.}$
Consider the following backward stochastic parabolic system with dynamic boundary conditions:
$\begin{cases}\begin{array}[]{ll}dy+\nabla\cdot(\mathcal{A}\nabla y)\,dt=(F_{1}+\partial_{t}h)\,dt+Y\,dW(t)&\text{in }Q,\\ dy_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}y\,dt=(F_{2}+\partial_{t}h_{\Gamma})\,dt+\widehat{Y}\,dW(t)&\text{on }\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\text{on }\Sigma,\\ (y(T),y_{\Gamma}(T))=(y_{T},y_{\Gamma,T})&\text{in }G\times\Gamma.\end{array}\end{cases}$ (2.5)
By Proposition 2.2, the backward system (2.5) admits a unique weak solution
$(y,y_{\Gamma};Y,\widehat{Y})\in\mathcal{Z}_{T}\equiv\Big(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\Big)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}),$
which satisfies, for any $t\in[0,T]$ and all $(\zeta,\zeta_{\Gamma})\in\mathbb{H}^{1}$,
$\begin{aligned} &\int_{G}(y(T)-y(t))\zeta\,dx+\int_{\Gamma}(y_{\Gamma,T}-y_{\Gamma}(t))\zeta_{\Gamma}\,d\sigma\\ &=\int_{t}^{T}\int_{G}\mathcal{A}\nabla y\cdot\nabla\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}(F_{1}+\partial_{s}h)\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}(F_{2}+\partial_{s}h_{\Gamma})\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}Y\zeta\,dx\,dW(s)+\int_{t}^{T}\int_{\Gamma}\widehat{Y}\zeta_{\Gamma}\,d\sigma\,dW(s),\quad\text{a.s.}\end{aligned}$ (2.6)
In particular, integrating by parts in time the third and fourth terms on the right-hand side of (2.6), we obtain the equivalent variational formulation
$\begin{aligned} &\int_{G}(y(T)-y(t))\zeta\,dx+\int_{\Gamma}(y_{\Gamma,T}-y_{\Gamma}(t))\zeta_{\Gamma}\,d\sigma\\ &=\int_{t}^{T}\int_{G}\mathcal{A}\nabla y\cdot\nabla\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}F_{1}\zeta\,dx\,ds-\int_{G}h(t)\zeta\,dx+\int_{t}^{T}\int_{\Gamma}F_{2}\zeta_{\Gamma}\,d\sigma\,ds-\int_{\Gamma}h_{\Gamma}(t)\zeta_{\Gamma}\,d\sigma\\ &\quad+\int_{t}^{T}\int_{G}Y\zeta\,dx\,dW(s)+\int_{t}^{T}\int_{\Gamma}\widehat{Y}\zeta_{\Gamma}\,d\sigma\,dW(s),\quad\text{a.s.}\end{aligned}$ (2.7)
Set
$(z,z_{\Gamma};Z,\widehat{Z})=(y,y_{\Gamma};Y,\widehat{Y})+(h,h_{\Gamma};0,0)\quad\text{in }(Q\times\Sigma)^{2},\;\text{a.s.}$